# DexSinGrasp: Learning a Unified Policy for Dexterous Object Singulation and Grasping in Cluttered Environments

### v1

@@ III-B2 Clutter Arrangement Curriculum Learning
Since it is challenging to learn object singulation and grasping with dexterous hands in compact or diverse environments, such as the D-8 or R-8 tasks, we design clutter arrangement curriculum to gradually increase the object diversity and spatial complexity. We begin by training a grasping policy designed exclusively for single-object scenarios, where the objective is to grasp a single block. Based on this initial policy, we continuously follow the curriculum and train on increasingly complex singulation and grasping tasks. Specifically, we first train on dense arrangements with D-4, then D-6, and finally D-8 tasks. We then use the expert obtained from the D-8 training as the starting point for training on random arrangements with R-4, R-6, and R-8 tasks. At the end of the training process, we extract the final policy from the D-8 training as the dense-clutter teacher policy (optimal on D-$n$ tasks) and the final policy from the R-8 training as the random-clutter teacher policy (optimal on R-$n$ tasks).
@@ III-C1 Data Collection for Distillation
The data collection phase involves preparing training data using two distinct teacher policies: the dense-clutter teacher policy for the D-4/6/8 tasks and the random-clutter teacher policy for the R-4/6/8 tasks, respectively. In total, 1000 episodes of observation and action data, along with scene point cloud, are prepared as $\{(s_{t}^{V},a_{t})\}$, where $s_{t}^{V}$ will be discussed later. The dataset is structured so that the D-4 and R-4 tasks each account for 10% of the total, while the D-6, D-8, R-6, and R-8 tasks each contribute 20%. This balanced distribution better captures cluttered environment tasks and ensures a comprehensive representation of varying task complexities, which is critical for effective policy distillation and better generalizability.
@@ III-C2 Vision-Based Policy Behavior Cloning
Our vision-based student policy uses scene point cloud instead of object poses, which cannot be accurately obtained in such heavily occluded, cluttered environments. We specifically use behavior cloning to train the student policy, with data collected from two teachers. The different visual observation $s_{t}^{V}$ for the vision-based student policy is defined as
@@
where singulation distance $d_{t}^{S}$ is removed from $s_{t}$ in the state-based teacher policy, and object state $s_{t}^{O}\in\mathbb{R}^{16}$ is substituted with center position of scene point cloud $s_{t}^{O^{\prime}}\in\mathbb{R}^{3}$ as the object state is hard to acquire. Moreover, the vision-based policy includes the visual features $v_{t}\in\mathbb{R}^{128}$ encoded from the scene point cloud using a pre-trained point cloud encoder from UniGraspTransformer [9], with the encoder weights frozen during the policy distillation.
@@ IV Experiment
Figure 4: Qualitative results on object singulation and grasping in simulation environments.
In this section, we conduct comprehensive experiments to evaluate our proposed method, DexSinGrasp, in both simulation and real-world tasks. Through these experiments, we aim to address the following key questions: (1) How effective and efficient is our method for grasping in clutter environments? (2) How does our method generalize to different objects and tasks? (3) How effective is our clutter arrangement curriculum learning? (4) How does our method perform on real-world tasks?
@@ IV-A Baselines
To evaluate our approach, we design three experimental configurations with two baseline methods and our proposed approach:
@@
Multi-Stage Singulation Method This baseline is a two-stage framework where separately trained singulation and grasping policies operate in sequence as adopted by SOPE [24]. We train the separate singulation policy without the grasping reward stage as mentioned in Sec. III-A. We also include a singulation bonus to encourage singulation. The singulation stage is switched to the grasping stage when $\sum_{i=1}^{n}\|p^{\text{target}}-p_{i}\|_{2}/n>0.16$, where $n$ is the number of surrounding objects.
@@ IV-B Evaluation Metrics
To comprehensively evaluate the performance of our approach, we introduce two key metrics that assess both effectiveness and efficiency in dexterous object singulation and grasping.
Success Rate. The proportion of trials in which the target object successfully reaches the predefined target position above the table surface is defined by $\|p^{\text{goal}}-p^{\text{target}}\|_{2}<0.05$. We denote the success rate as SR to evaluate the performance of each method.
Average Steps. The average number of simulation steps required to singulate and grasp the target object to goal positions. The unsuccessful trials are excluded from the calculation. We denote average steps as AS to evaluate the efficiency of each method.
@@ IV-C Implementation Details
We use Isaac Gym for clutter arrangement curriculum learning. For each D/R-$n$ task, we used 1000 simulated environments and trained the PPO policy network over 10K iterations with a learning rate of 3e-4. We then evaluated and selected the best-performing iteration as the policy for the next-stage clutter arrangement curriculum learning. The student policy is trained over 200 epochs with a batch size of 12 trajectories, each composed of 300 steps of recorded simulation data and a learning rate of 1e-4.
IV-D Main Results and Analysis
We tested the dense-clutter teacher policy and the distilled vision student policy on D-4, D-6, and D-8 tasks and compared their performance with the GraspReward-only and multi-stage singulation methods. All methods were evaluated in 10 environments over 10 episodes each, except the multi-stage singulation policy, which was tested in a single environment for 100 episodes due to its non-parallelizable stage-switching mechanism.
Based on the results presented in Tab. II, the multi-stage singulation policy demonstrates a higher success rate than the GraspReward-only baseline, suggesting that the singulation stage plays a positive role in task performance. Our dense-clutter teacher policy achieves a significantly higher average success rate of 98%, with a substantially lower AS compared to the multi-stage singulation policy. While the distilled vision student policy exhibits a lower SR than the teacher policy, it still outperforms the baseline policies. It maintains an AS comparable to that of the teacher policy. Our results show that while combining separate singulation and grasping policies can achieve a higher success rate, it increases action steps and reduces efficiency, whereas our unified policy balances effectiveness and efficiency for grasping in clutter environments, addressing Q1.
@@ IV-D Main Results and Analysis
GraspReward-Only 66% 40% 10% 39% 152 180 223 185
Multi-Stage Singulation 77% 76% 64% 72% 169 181 199 183
Ours (Teacher) 98% 99% 97% 98% 96 113 133 114
Ours (Student) 90% 92% 84% 89% 102 108 132 114
In response to Q2, we evaluate the random-clutter teacher policy and the distilled vision student policy on R-4, R-6, and R-8 tasks with different object arrangements to test the generalization ability of our policy. In Tab. III, we first observe the baseline policies perform better in random arrangements than dense arrangements, as the presence of relatively loose gaps provides more opportunities to grasp the target object directly. Our teacher policy can achieve an average SR of 96% across all tasks. While the distilled vision student policy exhibits an SR drop compared to the teacher policy, it still outperforms the baseline policies with a higher SR and lower AS for better task efficiency.
@@
GraspReward-Only 73% 61% 33% 56% 134 148 182 155
Multi-Stage Singulation 88% 72% 78% 79% 136 120 143 133
Ours (Teacher) 97% 96% 94% 96% 88 86 90 88
Ours (Student) 91% 86% 88% 88% 81 82 89 84
During training and testing, we found the policy learned several singulation patterns, including finger flickering, palm rubbing, and finger-palm vibration, to displace, nudge, or destabilize surrounding objects, effectively singulating targets in cluttered environments, as shown in Fig. 4.
@@ IV-E Clutter Arrangement Curriculum Learning Analysis
The clutter arrangement curriculum learning process is designed to enhance success rates in increasingly complex scenes. We evaluate each policy trained on D/R-$n$ tasks under various curriculum directions—dense to random (SO, D-4, D-6, D-8, R-4, R-6, R-8), random to dense (SO, R-4, R-6, R-8, D-4, D-6, D-8), and no curriculum (training each D/R-$n$ task from scratch)—as shown in Tab. IV. For the dense-to-random curriculum, we use the best-performing checkpoints at iterations 2k, 7.8k, 9.2k, 4.6k, 2.9k, 2.7k, and 2.5k respectively from trained 10k iterations; for the random-to-dense curriculum, at iterations 2k, 6.5k, 7.5k, 3.3k, 5.2k, 7.4k, and 1.6k respectively from trained 10k iterations. The curriculum with dense-to-random direction consistently yields the best performance across tasks. The results indicate that with the progression of the curriculum, the teacher policy demonstrate greater accuracy and efficiency, addressing Q3.
TABLE IV: Evaluation on different curriculums.
Curriculum SR(%$\uparrow$ on D-$n$ tasks SR(%$\uparrow$ on R-$n$ tasks
D-4 D-6 D-8 Avg. R-4 R-6 R-8 Avg.
Training from scratch 90% 75% 97% 87% 74% 26% 0% 33%
Random-to-dense 96% 87% 81% 88% 97% 92% 97% 95%
Dense-to-random 98% 92% 97% 96% 96% 96% 94% 95%
IV-F Real-World Experiments
Figure 5: Real-world experiment setting.
@@ V Conclusion
Our proposed approach demonstrates that a unified reinforcement learning framework can effectively integrate object singulation and grasping in densely or randomly cluttered environments using dexterous robotic hands. By unifying singulation and grasping, our method not only achieves higher grasp success rates and improved efficiency compared to traditional multi-stage approaches but also enables the dexterous hand to perform subtle non-prehensile maneuvers—such as nudging or sliding obstacles—directly during the grasping process. The integration of clutter arrangement curriculum learning and policy distillation further enhances the generalization of the vision-based policy, ensuring successful skill transfer from simulation to real-world applications. Additionally, the introduction of various cluttered grasping environments provides a comprehensive testbed for evaluating performance across various clutter configurations, reinforcing the superiority of our approach over conventional methods.
Future work can extend these promising results by addressing more complex object arrangements and incorporating a broader range of object shapes and dynamic clutter scenarios to push the limits of the current framework. Improving robustness in real-world settings by mitigating challenges such as sensor noise and sim-to-real discrepancies, as well as exploring alternative learning paradigms—including imitation learning and hierarchical control strategies—will be crucial for advancing dexterous manipulation. These directions offer valuable pathways toward developing a robust, general-purpose robotic system capable of efficient and reliable object retrieval in increasingly unstructured environments.

### v3

@@ III-B2 Clutter Arrangement Curriculum Learning
Since it is challenging to learn object singulation and grasping with dexterous hands in compact or diverse environments, such as the D-8 or R-8 tasks, we design clutter arrangement curriculum to gradually increase the object diversity and spatial complexity. We begin by training a grasping policy designed exclusively for Single-Object (SO) scenarios, where the objective is to grasp a single cuboid. Based on this initial policy, we continuously follow the curriculum and train on increasingly complex singulation and grasping tasks. Specifically, we first train on dense arrangements with D-4, then D-6, and finally D-8 tasks. We then use the expert obtained from the D-8 training as the starting point for training on random arrangements with R-4, R-6, and R-8 tasks. At the end of the training process, we extract the final policies from the D-8 training and the R-8 training as two teacher policies.
@@ III-C Teacher-Student Policy Distillation
Figure 3: Dense and random arrangement settings. We introduce a cluttered environment generation module to create diverse object settings.
@@ III-C1 Data Collection for Distillation
The data collection phase involves preparing training data using two distinct teacher policies: the dense-clutter teacher policy for the D-4/6/8 tasks and the random-clutter teacher policy for the R-4/6/8 tasks, respectively. In total, 1000 episodes of observation and action data, along with scene point cloud, are prepared as $\{(s_{t}^{V},a_{t})\}$, where $s_{t}^{V}$ will be discussed later. The dataset is structured so that the D-4 and R-4 tasks each account for 10% of the total, while the D-6, D-8, R-6, and R-8 tasks each contribute 20%. This balanced distribution better captures SNC tasks and ensures a comprehensive representation of varying task complexities, which is critical for effective policy distillation and better generalizability.
@@ III-C2 Vision-Based Policy Behavior Cloning
Our vision-based student policy uses scene point cloud instead of object poses, which cannot be accurately obtained in such heavily occluded SNCs. We specifically use behavior cloning to train the student policy, with data collected from two teachers. The different visual observation $s_{t}^{V}$ for the vision-based student policy is defined as
@@
where singulation distance $d_{t}^{S}$ is removed from $s_{t}$ in the state-based teacher policy, and object state $s_{t}^{O}\in\mathbb{R}^{16}$ is substituted with center position of scene point cloud $s_{t}^{O^{\prime}}\in\mathbb{R}^{3}$. Moreover, the vision-based policy includes the visual features $v_{t}\in\mathbb{R}^{128}$ encoded from the scene point cloud using a pre-trained point cloud encoder from UniGraspTransformer [9], with the encoder weights frozen during the policy distillation.
@@ IV Experiment
In this section, we conduct comprehensive experiments to evaluate our proposed method, DexSinGrasp, in both simulation and real-world tasks. Through these experiments, we aim to address the following key questions: (1) How effective and efficient is our method for grasping in SNCs? (2) How does our method generalize to different spatial arrangements? (3) How effective is our clutter arrangement curriculum learning? (4) How does our method generalize to different object geometries? (5) How does finger dexterity affect singulation and grasping performance in SNCs? (6) How does our method perform on real-world tasks?
@@ IV-A Baselines
To evaluate our approach, we compare our proposed approach with two baseline methods:
@@
Multi-Stage Method This baseline is a two-stage framework where separately trained singulation and grasping policies operate in sequence as adopted by SOPE [24]. We train the separate singulation policy without the grasping reward stage as mentioned in Sec. III-A. The singulation stage is switched to the grasping stage when $\sum_{i=1}^{n}\|p^{\text{target}}-p_{i}\|_{2}/n>0.16$, where $n$ is the number of surrounding objects.
@@ IV-B Evaluation Metrics
To evaluate the performance of our approach, we introduce two key metrics that assess both effectiveness and efficiency in dexterous object singulation and grasping.
Success Rate. The proportion of trials where the target object reaches the predefined goal position above the table, defined as $\|p^{\text{goal}}-p^{\text{target}}\|_{2}<0.05$. Denoted as SR.
Average Steps. The average number of steps to singulate and grasp the target object, denoted as AS. Failed episodes are counted as 300 steps (the maximum per episode).
The best results in each table are highlighted in bold. Parenthesized values show the difference from our method.
@@ IV-C Implementation Details
We use Isaac Gym for clutter arrangement curriculum learning. For each D/R-$n$ task, we used 1000 simulated environments and trained the PPO policy network over 10K iterations with a learning rate of $3\times 10^{-4}$. We then evaluated and selected the best-performing iteration as the policy for the next-stage clutter arrangement curriculum learning. The student policy is trained over 200 epochs with a batch size of 12 trajectories, each composed of 300 steps of recorded simulation data and a learning rate of $1\times 10^{-4}$. In both simulation and real-world experiments, hand rotation is disabled, restricting the end-effector to translation and finger motion. This encourages dexterous finger use for singulation and grasping, prevents unrealistic floating-hand movements that may cause arm–table collisions on the xArm, and eases sim-to-real transfer. We use Isaac Gym’s default friction (1.0) for experiments without domain randomization, and maintain small inter-object gaps (1–1.5 cm) to enable horizontal contact forces and avoid self-locking.
@@
State-Based GraspReward-Only 66 40 10 39 202 252 292 249
Multi-Stage 77 76 64 72 199 210 235 215
Ours (Teacher) 98 99 97 98 100 115 138 118
Vision-Based Ours (Student) 90 92 84 89 122 123 159 135
@@
State-Based GraspReward-Only 73 61 33 56 179 207 261 216
Multi-Stage 88 72 78 79 156 170 178 168
Ours (Teacher) 97 96 94 96 94 95 103 97
Vision-Based Ours (Student) 91 86 88 88 101 113 114 109
IV-D Main Results and Analysis
We tested the dense-clutter teacher policy and the student policy on D-4/6/8 tasks and compared their performance with all the baselines. All methods were evaluated for 100 times.
Based on the results presented in Tab. II, the multi-stage singulation policy demonstrates a higher success rate than the GraspReward-Only baseline, suggesting that the singulation stage plays a positive role in task performance. Our dense-clutter teacher policy achieves a significantly higher average success rate of 98%, with a substantially lower AS compared to the Multi-Stage policy. While the student policy has a lower SR than the teacher, it still outperforms the baselines and maintains a comparable AS. Our results show that combining separate singulation and grasping policies increases action steps and reduces efficiency, whereas our unified policy balances effectiveness and efficiency for grasping in cluttered environments including SNCs, thus addressing Q1.
In response to Q2, we evaluate the random-clutter teacher policy and the student policy on R-4/6/8 tasks with different object arrangements to test the spatial generalization ability of our policy. In Tab. III, we first observe that baselines perform better in random than in dense arrangements, where looser gaps allow direct grasps. Our teacher policy can achieve an average SR of 96% across all tasks. While the student policy achieves a lower SR than the teacher policy, it still outperforms all the baseliness in both effectiveness and efficiency.
During training and testing, we found the policy learned several singulation patterns, including finger flickering and finger-palm vibration to displace and nudge surrounding objects, effectively singulating targets from clutters, as shown in Fig. 4. These patterns rely heavily on the finger dexterity, and a more detailed discussion is provided in Sec. IV-G.
Figure 4: Qualitative results on object singulation and grasping in simulation.
TABLE IV: Evaluation on different curriculums.
[…]
D-4 D-6 D-8 Avg. R-4 R-6 R-8 Avg.
[…]
Dense-to-random (Ours) 98 92 97 96 96 96 94 95
@@ IV-E Clutter Arrangement Curriculum Learning Analysis
The clutter arrangement curriculum learning process is designed to enhance success rates in increasingly complex scenes. We evaluate each policy trained on D/R-$n$ tasks under various curriculum directions—dense to random (SO, D-4, D-6, D-8, R-4, R-6, R-8), random to dense (SO, R-4, R-6, R-8, D-4, D-6, D-8), and no curriculum (training each D/R-$n$ task from scratch)—as shown in Tab. IV. For the two curricula, we use the best-performing checkpoints from trained 10k iterations for each setting. The curriculum with dense-to-random direction consistently yields the best performance across tasks. The results indicate that with the progression of the curriculum, the teacher policy demonstrate greater accuracy and efficiency, addressing Q3.
IV-F Generalization to irregular clutters
To test whether our policy generalizes to geometries beyond the cuboid clutter, we evaluate its performance on randomly generated irregular clutters that are tightly packed—a particularly challenging case for SNCs. Each clutter is constructed by dividing a 24×24×8 cm cuboid using three random curves, ensuring three intersections to form seven distinct pieces, shown in Fig. 5. We fine-tuned our policy from D-8 expert on 200 such irregular clutters and tested it on 50 unseen instances together with previous baselines. As shown in Table V, our method achieves good performance, and therefore exhibits geometric generalization, in response to Q4.
Figure 5: Irregular Clutter. The target object is marked as purple in the center.
TABLE V: Evaluation on irregular clutters
@@ IV-F Generalization to irregular clutters
State-Based GraspReward-Only 26 271
Multi-Stage 53 245
Ours (Teacher, D-8 Finetuned) 81 211
Vision-Based Ours (Student, D-8 Finetuned) 60 231
TABLE VI: Evaluation of Teacher Policy with Varying Hand Dexterity.
[…]
Full-DoF LEAP Hand (Ours) 98 99 97 81
[…]
Full-DoF LEAP Hand (Ours) 100 115 138 211
IV-G Discussion of Dexterity
In singulation and grasping tasks for SNCs—where limited space restricts finger articulation—the use of high-degree-of-freedom (DoF) hands becomes essential. We evaluate three variants of the LEAP Hand with varying degrees of dexterity:
  • Low-DoF: The index, middle, and ring fingers are underactuated with fixed abduction–adduction, where each finger’s 3 joints are coupled into 1 DoF. The thumb retains 3 DOFs. (Total: 6 DoF)
  • Mid-DoF: Index, middle, and ring fingers have independently actuated 3 joints with fixed abduction–adduction. The thumb is fully actuated with 4 DoFs. (total: 13 DoF)
  • Full-DoF: A fully actuated LEAP Hand with control over all lateral and longitudinal joints of every digit. The thumb is fully actuated with 4 DoFs. (total: 16 DoF)
To quantify the impact of dexterity, we benchmarked these 3 embodiments across D-4/6/8 tasks and irregular clutters. The results, summarized in Table VI, show significant improvement with increased hand dexterity, in response to Q5. High finger dexterity enables complex contact interactions, which are critical in SNCs where efficient singulation requires delicate finger skills instead of coarse palm movement.
IV-H Real-World Experiments
Figure 6: Real-world experiments for D-8, R-8, irregular and practical clutters.
TABLE VII: Evaluation of Vision-Based Student Policy in Real-world.
Method D-8 R-8 Irregular Clutter Practical Clutter
GraspReward-Only 3/20 2/20 0/20 17/100
Multi-Stage 7/20 4/20 3/20 22/100
Ours 15/20 12/20 12/20 61/100
We conduct real-world experiments using a uFactory xArm6 robot equipped with a LEAP Hand and two side-view RealSense D435 RGB-D cameras, as shown in Fig. 6. The synchronized RGB-D data from both cameras are converted into point clouds, transformed to the world frame using calibrated extrinsics, and fused into a single scene point cloud, which is further refined using the Iterative Closest Point (ICP) algorithm. A predefined crop around the clutter center removes background noise, and the fused cloud is downsampled to 1,024 points before being fed into the vision-based policy.
We accomplish singulation and grasping for D-8, R-8, irregular clutter and practical clutter consisting of previously unseen practical objects randomly arranged in the real-world environment and conduct 20 trials for each setting (5 settings for practical clutter). To further improve sim-to-real transfer, we additionally retrained the teacher policy with domain randomization on hand joint damping, stiffness, and friction (hand friction varies from 0.7 to 1.8 and table friction varies from 0.4 to 1.3 in Isaac Gym). As shown in Tab. VII, the GraspReward-Only baseline struggles in cluttered environments due to the lack of singulation. In contrast, the Multi-Stage method suffers from the sensitivity to sim-to-real discrepancies, often causing failures when singulation generates unseen configurations. Notably, in practical clutter, our policy achieves a 61% success rate in a zero-shot setting, demonstrating strong generalization across both shape and spatial configurations, even though it was trained solely on a cuboid-shaped curriculum.
@@ V Conclusion
Our proposed approach demonstrates that a unified RL framework can effectively integrate object singulation and grasping in SNCs using dexterous hands. By unifying singulation and grasping, our method not only achieves higher grasp success rates and improved efficiency compared to traditional multi-stage approaches but also enables subtle maneuvers directly with finger dexterity. The combination of clutter arrangement curriculum learning and policy distillation further enhances the generalization of the vision-based policy, ensuring successful skill transfer from simulation to the real world, and from seen blocks to unseen objects. Additionally, the introduction of various cluttered grasping environments provides a comprehensive testbed for evaluating performance across various clutter configurations. Interestingly, a curriculum trained solely on cuboid objects exhibits strong zero-shot generalization to diverse practical clutters, suggesting that spatial relationships—rather than geometric variations alone—are critical to grasping in densely cluttered environments and validating our curriculum design.
Future work could extend to more complex clutters, broader object shapes, and improved real-world robustness, while further exploring how spatial diversity in clutter arrangement curricula enhance generalization to unseen clutters.